OpenStudy (anonymous):

1/y = 1/x + 1/7 Solve for X. I need help with this question.

OpenStudy (anonymous):

subtract \(\frac{1}{7}\) from both sides you get \[\frac{1}{y}-\frac{1}{7}=\frac{1}{x}\] now actually subtract, you get \[\frac{7-y}{7y}=\frac{1}{x}\] then take the reciprocal of both sides, i.e. flip them, you will get the answer

OpenStudy (anonymous):

what he did there was multiply 1/y by 7/7 and 1/7 by y/y then he subtracted